twice the sum of a number and 6 is the same as 3 more than the number. find the number
Answer by rfer(16322)

2(x+6)=x+3
2x+12=x+3
2x-x=3-12
x=-9
-------------------
2(-9+6)=-9+3
-18+12=-6
-6=-6